Please help to show work
absolute value 3x+1 =9

Answer by adamchapman(301)

3x+1 =9
3x = 9 -1
3x = 8
x = 8/3
I hope this helps.
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk
Adam